Ammara Manufacturing Compeny applies overheed on the hass of dinect leoour costs The compeny estimanes anranl overheed costs will be RMI,S labour costs will be 841,900000 Durisg the month of March, Amman on two jobs: AAl6l and B8171. The summary dala relatol to these two johs are ss ollons 20,000 and anaal deect Compony Manufacturing Costs locurred Purchased RM108,000 of raw monerials on accoust Factory labour of RMIS$2000 and RM3,000 employer pyoll tsxes Manufacoaring overicad of indirect materials and indeect labour foe RM119,600 Assignmeat of Costs Direct materials: Job AA161 RM54,000 and Job BB171 RM42,000 Direct labour. Job AAl6I RM104,000 and Job BBITI RMS2,000 Indirect lahosr: RM4,000 During the month, the company completed Job AA161 and sold it on account for RM300000. Jeb BB17I wis unly partially completed Required (a) Compute the prealetemived overheed re (2 marks) (13 monls) (3 marks) (b) Journslixe the March transactions. te) Determine the amount of under or over applied manufacturing overhead by using the manefacturing overhead T-account (d) Prepare journal estry so close aey balance in the manutiacturing overhend (2 marks) le cost of goods sold account. rroTAL 20 MARKS! Continued.
Answer #1

a)Predetermined overhead rate=total estimated overhead costs/estimated direct labour costs
=1520000/1900000
=0.8
b)Raw material(db)108000
Accounts payable(cr)108000

Factory labour(db)152000
Factory wages payable(Cr)144000
Employer payroll tax payable(cr)8000

manufacuring overhead(db)119600
indirect labour payable(cr)119600

c)AA161:
Applied overhead=0.8*104000=83200
BB171:
Applied overhead=0.8*52000=41600
Total overhead applied=83200+41600=124800
Actual overhead=119600
It is overapplied by 124800-119600=5200

d)Manufacturing overhead(db)5200
Cost of goods sold(Cr)5200